OpenStudy (anonymous):

Can someone help me? What the expression in factored form? x^2+13x+42

OpenStudy (anonymous):

ya sure i did this work yesterday xD

OpenStudy (tyteen4a03):

Surely you meant x^2+13x+42. ;)

OpenStudy (anonymous):

ok so first x two times right? and then 13+42 witch is easy and ur answer with the x :D

OpenStudy (tyteen4a03):

@Rere13 That is not correct, unless you want to get a factorized expression with complex numbers in it.

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

YA IT IS shes trying to solve the expression THINK

OpenStudy (anonymous):

I have a learning disability are you saying x^2=13+42 then 55x

OpenStudy (anonymous):

what learning disability?

OpenStudy (anonymous):

it takes a while for me to learn things...things that should be simple is very hard for me to learn..so the way you just put that is hard for me to understand..which is why I asked you that question

OpenStudy (lifeisadangerousgame):

You just want this to be factored?

OpenStudy (anonymous):

oh well im not like that and other people understand but since your ike that dont worry i wil got step by step with you :) not only to get this answer but to undertand all the other ones u have and for a test

OpenStudy (anonymous):

ok..and @LifeIsADangerousGame that is all it asks for

OpenStudy (anonymous):

what ok...? well i guess u dont want help bye

OpenStudy (lifeisadangerousgame):

Alright Mindless_girl I'm really bad at factoring as well haha! But let's give this a shot c:

OpenStudy (anonymous):

wait yes I want your help..I was saying ok as if im ready

OpenStudy (anonymous):

thanks @LifeIsADangerousGame

OpenStudy (lifeisadangerousgame):

let me get a paper lol

OpenStudy (lifeisadangerousgame):

It's x^2 + 13x + 42 right?

OpenStudy (lifeisadangerousgame):

@mindless_girl

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry my mom called me

OpenStudy (lifeisadangerousgame):

It's okay c: I like to start off with what I know. So I know I need to have this (x )(x ) I also know that everything is positive so I'm put to put a plus in both. (x + )(x + ) Now what times what will equal 42 but ALSO add up to 13?

OpenStudy (anonymous):

is it going to be a decimal?

OpenStudy (anonymous):

I know that 7*6 =42 and also add up to 13

OpenStudy (lifeisadangerousgame):

Right!:D So now we have those, and those are the two numbers we are going to put in our

OpenStudy (lifeisadangerousgame):

factor xD SO now, figure out which numbers go in which parentheses ^_^ And you're done!

OpenStudy (anonymous):

7 goes into x^2 and 6 goes into 13x

OpenStudy (lifeisadangerousgame):

like this? (x + 7)(x + 6)?

OpenStudy (anonymous):

yea

OpenStudy (lifeisadangerousgame):

Correct! You have just factored x^2 + 13x + 42 ^_^

OpenStudy (anonymous):

omg thank you. You are such a big help. You made it so easy. :)

OpenStudy (lifeisadangerousgame):

No problem c: Factoring is easy, just takes a little work ^_^